CHIRPED-PULSE FOURIER TRANSFORM MICROWAVE SPECTRUM OF THE LOW GLOBAL WARM-ING POTENTIAL REFRIGERANT ALTERNATIVE, 2,3,3,3-TETRAFLUOROPROPENE

\underline{MARK D. MARSHALL}, HELEN O. LEUNG, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000; JOHN S. MUENTER, Department of Chemistry, University of Rochester, Rochester, NY 14627.

A newly-constructed, chirped-pulse Fourier transform microwave spectrometer is used to obtain the microwave spectrum of 2,3,3,3-tetrafluoropropene, which has been proposed as an alternative to HFC-134a for use as a mobile air conditioning refrigerant. With a bandwidth of 6 GHz, this instrument has the potential to provide complete coverage of the microwave spectrum from 6 to 18 GHz in two steps, although smaller intervals may be utilized to increase the available power per unit bandwidth. Aided by ab initio calculations, the rotational spectrum of the normal isotopologue of 2,3,3,3tetrafluoropropene is assigned and analyzed to give rotational and centrifugal distortion constants. With moderate signal averaging, the signal-to-noise ratio of the spectrum indicates that it should be possible to observe singly substituted ^{13}C isotopologues in natural abundance.